OpenStudy (anonymous):

If g(x)=4x-7, find g(1/2) and g(a+2)

jimthompson5910 (jim_thompson5910):

g(x)=4x-7 g(1/2)=4(1/2)-7 Keep going and evaluate ... ------------------------------------------------------- g(x)=4x-7 g(a+2)=4(a+2)-7 Distribute then combine like terms to simplify

OpenStudy (anonymous):

TY

jimthompson5910 (jim_thompson5910):

yw